calculate the standard entropy change for the following reaction,
2Ag(2)O(s)-->4 AG(s)+o2(g)
given S*[Ag2O]=121.3 J/K*mol,S*[Ag(s)]=42.6J/k*mol, and S*[O29g0]=205.1J/K*mol
1
Expert's answer
2020-03-18T12:54:50-0400

The standard entropy change for the given reaction is


"\Delta S_0=\Sigma S_0(\text{products})-\Sigma S_0(\text{reactants})=\\
=(4\cdot42.6+205.1)-(2\cdot121.3)=132.9\space\frac{\text{J}}{\text{K}\cdot\text{mol}}."